# Mixture of Cluster-conditional LoRA Experts for Vision-language Instruction Tuning

### 08 Feb 2024 v3

@@ 3.4 Mixture of Cluster-conditional Experts
where $\text{top}_{k}(\cdot)$ keeps the $k$ largest entries unchanged and sets others to zero and $\mathbf{C}_{[\mathbf{x}]}$, which is shared among all the layers, denotes the learnable embedding of the cluster that $\boldsymbol{x}$ belongs to. This is the key for the model to choose proper task experts for the input data. To endow the clustering embedding with task information, we initialize it to be the centroid of the represented cluster. $\mathbf{W}_{\text{gate}}$ denotes trainable weights of the linear gate, $\boldsymbol{\epsilon}\sim N(0,\frac{1}{E})$ denotes a noise term that is designed to add some randomness to the expert choosing process, which encourages MoCLE to explore multiple combinations of experts during training. $\tau$ denotes a temperature hyperparamter. The output $\mathbf{y}_{i}$ is then computed as a weighted sum of the output of each expert on the input $\mathbf{x}_{i}$ as,
$\mathbf{y}_{i}=\sum_{e=1}^{E}G_{e}\mathbf{W}_{e}\mathbf{x}_{i},$ (2)
where $\mathbf{W}_{e}$ represents the linear projection weights of the $e$-th expert, and $G_{e}$, the $e$-th entry in $\mathbf{G}$, indicates the contribution of the $e$-th expert to $\mathbf{y}_{i}$.
In this formulation, the model learns to choose the expert that is suitable for the data by examining its cluster. In addition, as a task has different choices of expert at different layers, expert number scales well with the number of task clusters. We mitigate task conflicts with only a few experts.
@@ 3.5 Universal Expert
However, as shown in Sec. 4.4.1, we find that the formulation in Sec. 3.4 still hurts the generalization ability of the entire model. We hypothesize that this might be attributed to the fact that instruction-tuned models generalize to unseen tasks by training on many instructions, while in our case, each expert sees fewer instructions than the original dense model does. Hence, we propose a universal expert that could learn from all the training data. Specifically, we fix the number of activated experts to 1 (i.e., $k$ equals 1) and define $G_{\text{max}}$ as the maximum element in $\mathbf{G}$. Then the output for all the experts is expressed as,
$\mathbf{y}_{i}=\left(\sum_{e=1}^{E}G_{e}\mathbf{W}_{e}+(1-G_{\text{max}})\mathbf{W}_{u}\right)\mathbf{x}_{i}.$ (3)
In this formulation, we additionally train a universal expert parameterized by $\mathbf{W}_{u}$. Different from task experts that are activated only for specific model inputs, the universal expert is activated for any inputs. The final output is the weighted sum of outputs of one of the experts and the universal expert. The weighting for this universal expert is the complement of that of the chosen task expert. Consequently, the task expert learns distinct skills for certain tasks while the universal expert masters holistic understanding of the training corpus. The synergy between them offers both specialization and generalization for the LVLM.
@@ 4 Experiment
In this section, we empirically evaluate MoCLE on various downstream tasks in a zero-shot manner.
@@ 4.1 Settings and Implementation Details
In all the experiments, we utilize the InstructBLIP [11] model, particularly its Vicunna-7B variant [64], as the large vision-language model (LVLM) to assess the effectiveness of the proposed method. We encode the instructions of different datasets using the all-MiniLM-L6-v2 variant of the Sentence Transformer model [43] and cluster their embeddings into $64$ clusters via $k$-means clustering algorithm. During the vision-language instruction tuning, we initialize the model based on pre-trained InstructBLIP checkpoints [11] and apply the proposed MoCLE on the LLM (i.e., Vicunna-7B). Specifically, for each transformer layer in the LLM, we insert the MoCLE, which consists of 4 task LoRA experts and one LoRA universal expert, into the query and value projection of self-attention blocks. Inside each expert, we set the rank of the LoRA to 8 and the temperature $\tau$ to $0.05$. During instruction tuning, we only finetune parameters of Q-Former and the LoRA experts while keeping other parts of the LLM frozen.
Flickr 30K GQA VSR IconQA TextVQA Visdial HM VizWiz SciQA image MSVD QA MSRVTT QA iVQA Trained using private data Flamingo-80B 67.2 - - - 35.0 - 46.4 31.6 - 35.6 17.4 40.7 Qwen-VL (7b) 85.8 $*{}^{*}$ $*{}^{*}$ 35.2 67.1 - - - Qwen-VL (7b)-chat 81.0 $*{}^{*}$ $*{}^{*}$ 38.9 68.2 - - - Trained using grounding data MiniGPT-v2 (7B) - $*{}^{*}$ 60.6 47.7 - - 58.2 32.9 - - - - MiniGPT-v2 (7B)-chat - $*{}^{*}$ 62.9 51.5 - - 58.8 53.6 - - - - Trained using public data LLaVA (13B) - 41.3 51.2 43.0 - - - - - - - - MiniGPT-4 (13B) - 30.8 41.6 37.6 - - - - - - - - BLIP-2 (7B) 74.9 38.6 50.0 39.7 40.1 44.9 50.6 25.3 53.8 18.3 9.2 27.5 InstructBLIP (7B) 82.4 49.2 54.3 43.1 50.1 45.2 59.6 34.5 60.5 41.8 22.1 52.2 Under the same evaluation InstructBLIP (${}^{{\dagger}}$ 82.9 49.2 65.2 43.1 50.9 46.0 64.3 33.7 59.4 41.5 22.7 51.7 InstructBLIP (${}^{{\ddagger}}$ 81.3 48.6 60.8 43.4 53.9 46.3 65.3 34.5 62.0 41.4 23.0 51.3 InstructBLIP (${}^{{\ddagger}}$+MoCLE 81.9 (+0.6) 49.3(+0.7) 64.7(+3.9) 46.3(+2.9) 57.1(+3.2) 46.9(+0.6) 65.6(+0.3) 34.5(+0.0) 63.9(+1.9) 42.6(+1.2) 24.4(+1.4) 53.2(+1.9)
Table 1: Zero-shot results on the held-out datasets. Here, Visdial, HM and SciQA denote Visual Dialog, HatefulMemes and ScienceQA, respectively. For ScienceQA, we only evaluate the set with image context. Following previous works [11], we report the CIDEr score [52] for Flickr30K, accuracy for iVQA, AUC score for HatefulMemes, and Mean Reciprocal Rank (MRR) for Visual Dialog. For all other datasets, we report top-1 accuracy (%). Notice results with * means the training split of the evaluation dataset is used during instruction tuning, thus not zero-shot. Best results obtained by models trained with public data are marked in bold and the second best are underlined. Superscript ${}^{{\dagger}}$ denotes the official InstructBLIP checkpoint. Superscript ${}^{{\ddagger}}$ denotes our reproduced version, into which MoCLE is incorporated.
@@ 4.2 Datasets and Instructions
For evaluation datasets, we follow InstructBLIP but only keep Flickr30K [58], GQA [19], VSR [32], IconQA [37], TextVQA [50], Hateful Memes [22], VizWiz [16], ScienceQA [38], MSVD-QA [56], MSRVTT-QA [56] and iVQA [57]. InstructBLIP originally uses the NoCaps dataset [1] for the image captioning task, which is not adopted here since we cannot reproduce the reported results, probably due to the incomplete Web CapFilt dataset used.
The instructions to conduct all the experiments in this paper are from the original InstructBLIP paper. The detailed introductions of the held-in and held-out tasks/datasets are also presented in [11].
4.3 Zero-shot Evaluations
Table 1 presents the experimental results on multiple zero-shot vision-language tasks. We build our method upon the Vicunna-7B variant of InstructBLIP. The InstructBLIP (${}^{{\ddagger}}$ is our reproduced version of InstructBLIP. Compared to this model, the proposed MoCLE method shows considerable performance improvement. Specifically, on conventional tasks such as visual question answering (VQA), we observe a 2.9% improvement on IconQA. On tasks with novel formats such as Visual Spatial Reasoning (VSR), TextVQA and ScienceQA image, we obtain an absolute performance gain of 3.9%, 3.2%, and 1.9%, respectively. This indicates that the proposed MoCLE method facilitates generalization to unseen tasks/instructions since the instructions of VSR, TextVQA and ScienceQA image are unseen during training. In addition, the proposed MoCLE method performs well when compared with existing popular LVLMs including LLaVA (13B) and MiniGPT-4 (13B). For example, on IconQA, the proposed MoCLE method outperforms LLaVA (13B) and MiniGPT-4 (13B) by 3.3% and 8.7%, respectively. In summary, the proposed MoCLE model obtains best results on IconQA, TextVQA, HM, and ScienceQA image across models trained using public data.
@@ 4.4 Ablation Studies
In this section, we first ablate the effectiveness of the main components (e.g., task experts and universal expert) in the proposed MoCLE method in Sec. 4.4.1. Then in Sec. 4.4.2 we conduct a thorough analysis to study how the proposed MoCLE method would respond to the change of hyper-parameters (e.g., temperature and the number of clusters and task experts). Notice that for the ablations, we use Flickr30K (Flickr), Hateful Memes (HM), ScienceQA (SciQA), IconQA, Visual Spatial Reasoning (VSR) and TextVQA for evaluation.
Components Flickr HM SciQA IconQA VSR TextVQA
InstructBLIP 81.3 65.3 62.0 43.4 60.8 53.9
+LoRA (rank=8) 81.3 65.1 57.4 44.2 62.8 49.4
+LoRA (rank=64) 81.5 65.2 62.0 43.9 62.6 49.0
+Mixture (Cluster) 81.9 65.4 63.3 46.1 58.9 54.9
+Universal Expert 81.9 65.6 63.9 46.3 64.7 57.1
@@
Table 2: Comparison of individual components of the MoCLE framework in zero-shot vision-language tasks. We include three ablation studies that cumulatively add (i) parameter-efficient finetuning of LLM using LoRA of different ranks, (ii) replacing the single LoRA with its mixture and route inputs by their cluster, and (iii) training with the universal expert. Default settings are marked in gray.
4.4.1 Effects of Different Components
@@
We start from vanilla InstructBLIP (Vicunna-7B) and add the main components in the proposed MoCLE method step by step. As shown in Table 2, naively applying LoRA to InstructBLIP could be problematic. Although we observe minor performance gain in IconQA and VSR, the peformance on ScienceQA and TextVQA are much inferior to InstructBLIP. We speculate that a single LoRA offers very limited flexibility, thus it performs not so well on many tasks. Note that we further conduct an experiment with a rank of 64 (much larger than $r=8$ in the default setting of MoCLE). However, the resulting model is still outperformed by InstructBLIP with less trainable parameters. This indicates that naively increasing the capacity of the network cannot mitigate the task conflicts in the training data. However, when we replace a single LoRA with the Mixture of LoRAs and route the input based on the cluster of their instructions, the resulting model outperforms its predecessor by a large margin on SciQA, IconQA, and TextVQA. Nevertheless, we notice a sharp performance drop in the VSR task. This could be due to the fact that instruction-tuned model generalizes to unseen tasks by training on many instructions, while in our case, each expert sees fewer instructions than the original dense model does. As shown in Table 2, this issue could be alleviated by the use of the universal expert, which is updated by the gradients from all the training data. The same result can be observed in TextVQA, which requires not only the VQA ability but also additional optical character recognition (OCR) skills. Specifically, in the training stage, we have instructions “Question: {}. Short answer:” for all the VQA datasets and “OCR tokens: {}. Write a description for the photo” for TextCaps [49], but the instruction of TextVQA [50] is “OCR tokens: {}, Question: {}. Short answer:”. This requires the model to combine the knowledge from two tasks.
Universal Expert vs. Top-2 Experts.
@@
MoCLE shares similarities with the top-2 variant of MoE. Specifically, they both activate two experts each time. However, the universal expert differs from top-2 MoE by consistently selecting the universal expert, whose contribution is the complement of the top-1 expert. As shown in Table 3, the top-2 MoE model yields inferior results compared to MoCLE. Therefore the naively activating one more expert cannot address the generalization problem. Surprisingly, this variant performs even worse than the MoCLE model without the universal expert reported in Table 2. This might be resulted from task conflicts with more parameters shared.
Flickr HM SciQA IconQA VSR TextVQA
Universal 81.9 65.6 63.9 46.3 64.7 57.1
top-2 82.0 64.7 61.9 45.5 56.3 52.0
@@
Gating strategy Flickr HM SciQA IconQA VSR TextVQA
Token 81.7 65.4 61.9 44.0 49.0 46.6
Sentence 82.0 65.1 62.3 45.3 56.6 47.0
Cluster 81.9 65.6 63.9 46.3 64.7 57.1
@@
Gating Strategies.
We compare the proposed cluster-conditioned gating strategy with two popular ones: Token-MoLE and Sentence-MoLE, where MoLE denotes the mixture of LoRA experts by applying the MoE technique to LoRA, Token-MoLE obtains the routing decision based on the hidden representations of each token, and Sentence-MoLE averages the representations of the instruction tokens while excluding the visual tokens. As observed in Table 4, Token-MoLE and Sentence-MoLE give inferior results on the evaluation tasks. We speculate this is because that (1) an expert only learns a subset of the tasks, giving poor generalizations and (2) similar tasks are not grouped together by the same expert, resulting in task conflicts within that expert. The later can be evidenced by the routing visualization in Figure 5, where samples in the same dataset are routed to multiple experts instead of a dedicate one.
4.4.2 Hyper-parameter Analysis
In this section, we analyze in details the effect of the hyper-parameter in the MoCLE framework to the performance. Specifically, we investigate the how would MoCLE react to the number of clusters as introduced in Sec. 3.3 and the number of task experts as well as the temperature value in Sec. 3.4. If not otherwise specified, we employ all the components ablated in Sec. 4.4.1 and use a default configuration of 64 clusters with 4 experts and a temperature of 0.05.
@@ 4.4.2 Hyper-parameter Analysis
Temperature $\tau$ Flickr HM SciQA IconQA VSR TextVQA
0.01 81.6 64.7 61.7 46.3 61.2 55.8
0.05 81.9 65.6 63.9 46.3 64.7 57.1
0.1 82.3 65.7 63.7 45.7 63.0 56.7
0.2 81.9 64.8 62.1 45.2 61.3 54.0
Table 7: Ablation study on the temperature $\tau$ in Eq. (1).
Number of Clusters.
@@
Number of Task Experts.
@@
Figure 3: t-SNE visualization of the instruction encoding. Different colors correspond to different datasets, while the shape of the markers indicates the task category defined manually.
Temperature.

### 22 Mar 2024 v4

@@ 3.4 Mixture of Cluster-conditional LoRA Experts
where $\text{top}_{k}(\cdot)$ keeps the $k$ largest entries unchanged and sets others to zero and $\mathbf{C}_{[\mathbf{x}]}$, which is shared among all the layers, denotes the learnable embedding of the cluster that $\boldsymbol{x}$ belongs to. This is the key for the model to choose proper task experts for the input data. To endow the clustering embedding with task information, we initialize it to be the centroid of the corresponding cluster. $\mathbf{W}_{\text{gate}}$ denotes trainable weights of the linear gate, $\boldsymbol{\epsilon}\sim N(0,\frac{1}{E})$ denotes a noise term that is designed to add some randomness to the expert choosing process, which encourages MoCLE to explore multiple combinations of experts during training. $\tau$ denotes a temperature hyperparamter. The output $\mathbf{y}_{i}$ is then computed as a sum of the weighted output of each expert and the original LLM linear layer [21] on the input $\mathbf{x}_{i}$ as,
$\mathbf{y}_{i}=\sum_{e=1}^{E}G_{e}\mathbf{W}_{e}\mathbf{x}_{i}+\mathbf{W}_{0}\mathbf{x}_{i},$ (2)
where $\mathbf{W}_{0}$ denotes the pre-trained linear layer of LVLMs, $\mathbf{W}_{e}$ is the linear projection weights of the $e$-th LoRA expert, and $G_{e}$, the $e$-th entry in $\mathbf{G}$, indicates the contribution of the $e$-th expert. In this formulation, the model learns to choose the expert that is suitable for the data by examining its cluster. We will show in Sec. 4.4.3 that the condition of clusters in expert routing is indispensable for task generalization with significant superiority over typical routing conditions including sentence embeddings and token representations.
@@ 3.5 Universal Expert
However, as shown in Sec. 4.4.1, we find that the formulation in Sec. 3.4 still hurts the generalization ability of the entire model. We hypothesize that this might be attributed to the fact that the instruction-tuned models generalize to unseen tasks by training on extensive instructions [60], while in our formulation, each expert sees fewer instructions than the original dense model. Hence, we propose a universal expert that could learn from all the training data. Specifically, we fix the number of activated experts to 1 (i.e., $k$ equals 1) and define $G_{\text{max}}$ as the maximum element in $\mathbf{G}$. Then the output for all the experts is expressed as,
$\mathbf{y}_{i}=\left(\sum_{e=1}^{E}G_{e}\mathbf{W}_{e}+(1-G_{\text{max}})\mathbf{W}_{u}\right)\mathbf{x}_{i}+\mathbf{W}_{0}\mathbf{x}_{i}.$ (3)
In this formulation, we additionally train a universal expert parameterized by $\mathbf{W}_{u}$. Different from task experts that are activated only for specific model inputs, the universal expert is activated for any inputs. The final output is the weighted sum of outputs from one of the experts and the universal expert puls the original LVLM’s output. The weighting for this universal expert is the complement of that of the chosen task expert. Consequently, the task expert learns distinct skills for certain tasks while the universal expert masters holistic understanding of the training corpus. The synergy between them offers both specialization and generalization for the LVLMs with MoCLE.
@@ 4 Experiment
In this section, we conduct an assessment of MoCLE across multiple downstream tasks under a zero-shot setting. Initially, we detail the experimental settings and implementation details, which is followed by a description of the datasets and instructions employed, along with the outcomes of our evaluations. Lastly, we present an ablation study and visualizations of clustering and routing results.
@@ 4.1 Settings and Implementation Details
In all the experiments, we utilize the InstructBLIP[14] model, particularly its Vicuna-7B variant [70], as the base LVLM to assess the effectiveness of MoCLE. We encode all the instructions of different datasets using the all-MiniLM-L6-v2 variant of the Sentence Transformer model [49] and cluster their embeddings into $64$ clusters via $k$-means clustering algorithm. We apply the proposed MoCLE on the LLM (i.e., Vicuna-7B) of InstructBLIP. Specifically, for each transformer layer in the LLM, we insert the MoCLE, which consists of 4 task LoRA experts and one LoRA universal expert, into the query and value projection of self-attention blocks. Inside each expert, we set the rank of the LoRA to 8 and the temperature $\tau$ to $0.05$. During instruction tuning, we only finetune parameters of Q-Former and the LoRA experts while keeping other parts of the LLM frozen.
Models GQA VSR IconQA Visdial MME POPE A-OKVQA OKVQA VQAv2 Direct MC (test) (test-dev) LLaVA (13B) 41.3 51.2 43.0 - - 67.0 50.5 - 54.4 - MiniGPT-4 (13B) 30.8 41.6 37.6 - 581.7 71.4 38.7 - 37.5 - InstructBLIP (7B) 48.6 60.8 43.4 46.3 1202.9 77.6 58.8 73.8 57.0 77.4 49.3 64.7 46.3 46.9 1222.6 82.1 61.5 78.2 59.8 78.9 MoCLE (7B) (+0.7) (+3.9) (+2.9) (+0.6) (+19.7) (+4.5) (+2.7) (+4.4) (+2.8) (+1.5)
Table 1: Zero-shot results on the held-out datasets (GQA, VSR, IconQA, Visdial) and LVLM benchmarks (MME, POPE) and evaluation on held-in datasets (A-OKVQA, OKVQA, VQAv2). Here Direct and MC denote directly answering and multiple choices. Best results are marked in bold.
Models Flickr 30K TextVQA HM SciQA image MSVD QA MSRVTT QA iVQA InstructBLIP (7B) 81.3 53.9 65.3 62.0 41.4 23.0 51.3 81.9 57.1 65.6 63.9 42.6 24.4 53.2 MoCLE (7B) (+0.6) (+3.2) (+0.3) (+1.9) (+1.2) (+1.4) (+1.9)
Table 2: Zero-shot results on the held-out datasets. Here, HM and SciQA denote HatefulMemes and ScienceQA, respectively.
@@ 4.2 Datasets and Instructions
For the zero-shot evaluation datasets, we follow InstructBLIP but only keep Flickr30K [64], GQA [22], VSR [36], IconQA [43], TextVQA [56], Hateful Memes [25], ScienceQA [42], MSVD-QA [61], MSRVTT-QA [61] and iVQA [62]. NoCaps dataset [1] is originally utilized for the image captioning task, which, however, is not adopted here since we cannot reproduce the reported results, probably due to the incomplete Web CapFilt dataset used. In addition, we consider another two widely adopted LVLM benchmarks, i.e., MME [16] and POPE [33] for zero-shot evaluations. For held-in evaluation, we consider the VQAv2 [19], OKVQA [45] and A-OKVQA [52].
For ScienceQA, we only evaluate the set with image context. Following previous works [14], we report the CIDEr score [57] for Flickr30K, the iVQA accuracy for iVQA, AUC score for HatefulMemes, and Mean Reciprocal Rank (MRR) for Visual Dialog. For LVLM benchmarks, we report the perception score for MME and F1 score for the adversarial split of POPE. For all other datasets, we report the top-1 accuracy (%). The instructions to conduct all the experiments in this paper are from the original InstructBLIP paper. The detailed introductions of the held-in and held-out tasks/datasets are also presented in [14].
4.3 Evaluation Results
Table 1 and Table 2 present the experimental results on multiple zero-shot and held-in vision-language tasks. We build our method upon the Vicuna-7B variant of InstructBLIP. Compared to this model, the proposed MoCLE method shows considerable performance improvement. Specifically, on held-out datasets such as IconQA, Visual Spatial Reasoning (VSR), TextVQA and ScienceQA datasets, we obtain an absolute performance gain of 2.9% 3.9%, 3.2%, and 1.9%, respectively. This indicates that the proposed MoCLE method facilitates generalization to unseen tasks/instructions since these datasets are unseen during training.
In addition, we observe considerable improvement on common LVLM benchmarks (i.e., MME and POPE). POPE evaluates whether the model suffers from object hallucinations while MME considers a wide range of abilities such as math, OCR, color and understanding of arts. Therefore, our MoCLE cannot only promote task generation, but also improve overall abilities of the LVLM.
Further, the held-in tasks whose datasets are seen during training could also benefit from the proposed MoCLE. Specifically, with our MoCLE, an absolute improvement of 4.4%, 2.8% and 1.5% could be observed on A-OKVQA (MC), OKVQA and VQAv2. This indicates that MoCLE could effectively alleviate the task conflicts during multi-task learning.
@@ 4.4 Ablation Studies
In this section, we first ablate the effectiveness of the main components (e.g., task experts and universal expert) in the proposed MoCLE method. Then we conduct a thorough analysis to study how the proposed MoCLE method would respond to the change of hyper-parameters (e.g., temperature and the number of clusters and task experts). Notice that for the ablations, we use Flickr30K (Flickr), Hateful Memes (HM), ScienceQA (SciQA), IconQA, Visual Spatial Reasoning (VSR) and TextVQA for evaluation.
Components Flickr HM SciQA IconQA VSR TextVQA Avg.
InstructBLIP 81.3 65.3 62.0 43.4 60.8 53.9 61.1
+LoRA (rank=8) 81.3 65.1 57.4 44.2 62.8 49.4 60.0
 (rank=64) 81.5 65.2 62.0 43.9 62.6 49.0 60.7
+Mixture (Cluster) 81.9 65.4 63.3 46.1 58.9 54.9 61.8
+Universal Expert 81.9 65.6 63.9 46.3 64.7 57.1 63.3
@@
Table 3: Comparison of individual components of the MoCLE framework in zero-shot vision-language tasks. We include three ablation studies that cumulatively add (i) parameter-efficient finetuning of LLM using LoRA of different ranks (higher ranks are ablated, i.e. rank=64, but we use rank=8 for ablations and MoCLE), (ii) replacing the single LoRA with its mixture and route inputs by their cluster, and (iii) training with the universal expert. Default settings are marked in gray.
4.4.1 Effects of Different Components.
@@
We start from vanilla InstructBLIP with Vicuna-7B and add the main components in the proposed MoCLE method step by step. As shown in Table 3, naïvely applying LoRA to InstructBLIP could be problematic. Although we observe minor performance gain in IconQA and VSR, the peformance on ScienceQA and TextVQA are much inferior to InstructBLIP. We speculate that a single LoRA offers very limited flexibility, thus it performs not so well on many tasks. Note that we further conduct an experiment with a rank of 64 (much larger than $r=8$ in the default setting of MoCLE). However, the resulting model is still outperformed by InstructBLIP with less trainable parameters. This indicates that naïvely increasing the capacity of the network cannot mitigate the task conflicts in the training data. However, when we replace a single LoRA with the Mixture of LoRAs and route the input based on the cluster of their instructions, the resulting model outperforms its predecessor by a large margin on SciQA, IconQA, and TextVQA. Nevertheless, we notice a sharp performance drop in the VSR task. This could be due to the fact that instruction-tuned model generalizes to unseen tasks by training on many instructions, while in our case, each expert sees fewer instructions than the original dense model does. As shown in Table 3, this issue could be alleviated by the usage of the universal expert, which is updated by the gradients from all the training data. The same result can be observed in TextVQA, which requires not only the VQA ability but also additional optical character recognition (OCR) skills. Specifically, in the training stage, we have instructions “Question: {}. Short answer:” for all VQA datasets and “OCR tokens: {}. Write a description for the photo” for TextCaps [55], while the TextVQA [56] instruction is “OCR tokens: {}, Question: {}. Short answer:”, requiring LVLMs to combine knowledge from both tasks.
4.4.2 Universal Expert vs. Top-2 Experts.
@@
MoCLE shares similarities with the top-2 variant of MoE. Specifically, they both activate two experts each time. However, the universal expert differs from top-2 MoE by consistently selecting the universal expert, whose contribution is the complement of the top-1 expert. As shown in Table 4, the top-2 MoE model yields inferior results compared to MoCLE. Therefore the naïvely activating one more expert cannot address the generalization problem. Surprisingly, this variant performs even worse than the MoCLE model without the universal expert reported in Table 3. This might be resulted from task conflicts with more parameters shared.
Flickr HM SciQA IconQA VSR TextVQA Avg.
Universal 81.9 65.6 63.9 46.3 64.7 57.1 63.3
Top-2 82.0 64.7 61.9 45.5 56.3 52.0 60.4
@@
Gating strategy Flickr HM SciQA IconQA VSR TextVQA Avg.
Token 81.7 65.4 61.9 44.0 49.0 46.6 58.1
Sentence 82.0 65.1 62.3 45.3 56.6 47.0 59.7
Dataset 80.3 64.6 63.1 45.9 62.6 55.8 62.1
Cluster 81.9 65.6 63.9 46.3 64.7 57.1 63.3
@@
4.4.3 Gating Strategies.
We compare the proposed cluster-conditioned gating strategy with Token-MoLE, Sentence-MoLE and Dataset-MoLE, where MoLE denotes the mixture of LoRA experts by applying the MoE technique to LoRA, Token-MoLE obtains the routing decision based on the hidden representations of each token, and Sentence-MoLE averages the representations of the instruction tokens while excluding the visual tokens. Dataset-MoLE is a special case of MoCLE as it treats each dataset as a cluster while MoCLE leverages $k$-means to obtain clusters among all the datasets. Note that the universal expert is adopted for all these MoLEs. As observed in Table 5, Token-MoLE and Sentence-MoLE give inferior results on the evaluation tasks. We speculate that this is because (1) an expert only learns a subset of the tasks, giving poor generalization and (2) similar tasks are not grouped together by the same expert, resulting in task conflicts within that expert. The later can be evidenced by the routing visualization in Figure 4, where samples in the same dataset are routed to multiple experts instead of a dedicate one. Further, we observe performance drops when replacing $k$-means clustering with statically treating each dataset as a cluster. This results from the fact that Dataset-MoLE is less flexible than MoCLE because it can only assign a dataset to one cluster. However, in practice, we observe multiple tasks in a dataset which should be assigned to different clusters. (e.g., llava_150k contains reasoning, conversations and captioning, which are assigned to different clusters/experts as in Figure 2(a) and 3(a)).
Table 6: Ablation study on the number of clusters $K$.
@@
4.4.4 Number of Clusters.
@@
4.4.5 Number of Task Experts.
@@
4.4.6 Temperature.